Suppose that Hubert and Kate are the only suppliers of ice cream cones in a particular market. The following table shows their monthly supply schedules Huberts Quantity Supplied (Cones) Kates Quantity Supplied (Cones) Price (Dollars per cone) 13 14 On the following graph, plot Huberts supply of ice cream cones using the green points (triangle symbol). Next, plot Kates supply of ice cream cones using the purple points (diamond symbol). Finally, plot the market supply of ice cream cones using the orange points (square symbol) Huberts Supply Kates Supply Market Supply 12 16 20 24 QUANTITY (Cones)
Answer #1

Market Supply Curve is the curve which shows all the combinations of Sum total of quantity supplied by all individual firm at a particular price.

Market Supply Schedule.

Price    Quantity Supplied

1    0+4=4

2 4+8=12

3    6+11=17

4    7+13= 20

5    8+14 =22

# A point to make is that Supply curve is not necessarily linear. Though It is always upward sloping.

Maskut Supp Hubos Subtl Kab Subb 12. 16 28

ALL THE BEST :))